OpenStudy (anonymous):

The dimensions of a smaller rectangle are 3 ft. by 9 ft. The dimensions of a larger rectangle are 5 ft. by 11 ft. Find the ratio of the perimeter of the larger rectangle to the perimeter of the smaller rectangle. Options are 11:9, 3:4, 4:3, and 5:3

OpenStudy (anonymous):

perimeter of smaller rect = 2L + 2B = 2*3ft + 2*9ft = 2(3 + 9) ft perimeter of bigger rect = 2L + 2B = 2*5ft + 2*11ft = 2(5 + 11) ft perimeter of larger rect 2(3 + 9) ft 12 3 --------------------- = ----------- = ------ = ----- perimeter of smaller rect 2(5 + 11) ft 16 4 so their ratio is 3:4

OpenStudy (mckenzieandjesus):

@jim_thompson5910 was the person above right or is is 4:3?